OpenStudy (anonymous):

Simplify. 5^-1(3^-2) explain please? a. 45 b. 1/45 c. 15^-3 d. 15^2

OpenStudy (compassionate):

\(\Huge\bf Welcome~to~ \color{#00B4ff}{Open}\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge\bf\color{#7cc517}{Study~!!}~\) Can you write the expression using the Equation button at the bottom left?

OpenStudy (usukidoll):

negative exponents aren't allowed...therefore they are written as 1 / whatever the number is so .. for 5^-1 we have 1/5^1 which is 1/5. for 3^-2 we hace 1/3^2 which is 1/9 1/5 multiplied by 1/9 is 1/45

OpenStudy (usukidoll):

\[5^{-1}\] \[\frac{ 1 }{ 5^1 }\] \[\frac{ 1 }{ 5 }\]

OpenStudy (usukidoll):

\[3^{-2}\] \[\frac{ 1 }{ 3^2}\] \[\frac{ 1 }{ 9}\]

OpenStudy (usukidoll):

\[\frac{ 1 }{ 9}\] X \[\frac{ 1 }{ 5}\] = \[\frac{ 1 }{ 45}\]

OpenStudy (anonymous):

Thank you so much!!! That helped a lot :)